NAME (print neatiy) Score (last name, first name) Markup Problem. Consider a situation, where a machine shop is bidding a contract for to provide an aluminum manifold. The costs per unit are $51.36 and the price the component will sell for is $96.42. Important to show all work and provide appropriate units with two decimal places. 1. What is the markup based on cost and price? (1 point for each part Markup as a percentage of cost Markup as a percentage of price 2. How many units must be sold to achieve a profit of S175k? 2 points) 3. Now, assume costs of $57.51, what would the unit price need to be if the manufacturer 4. Now, assume a price of $91.77, what would the unit costs need to be if the manufacturer used a markup as a percentage of the cost of 39%? (3 points)

PLEASE SHOW WORK OF QUESTIONS 1-4
Answer #1
Particulars Amount
Sell price per unit 96.42
Cost per unit 51.36
Profit per unit 45.06
Markup based on selling price= Profit per unit / selling price
Markup based on selling price= 45.06/96.42 i.e 46.73%
Markup based on cost = Profit per unit / Cost price per unit
Markup based on cost = 45.06/51.36 i.e 87.73%
2) Desired profit = 175000
Desired profit = Unit sold * profit per unit
175000 = Units sold * 45.06 per unit
Units sold = 175000/ 45.06 i.e 3883.71
3) Cost = 57.51
Let selling price be x
X - 57.51 = 0.39*X
X- 57.51 = 0.39X
0.61 X = 57.51
X = 57.51/0.61
X= 94.28
Selling price is to be 94.28
4) Selling price = 91.77
Let unit cost be x
X+ 0.39X = 91.77
1.39X = 91.77
X = 66.02
Cost is 66.02